### v1

# MEGADance: Mixture-of-Experts Architecture for Genre-Aware 3D Dance Generation

Abstract
Music-driven 3D dance generation has attracted increasing attention in recent years, with promising applications in choreography, virtual reality, and creative content creation. Previous research has generated promising realistic dance movement from audio signals. However, traditional methods underutilize genre conditioning, often treating it as auxiliary modifiers rather than core semantic drivers. This oversight compromises music-motion synchronization and disrupts dance genre continuity, particularly during complex rhythmic transitions, thereby leading to visually unsatisfactory effects. To address the challenge, we propose MEGADance, a novel architecture for music-driven 3D dance generation. By decoupling choreographic consistency into dance generality and genre specificity, MEGADance demonstrates significant dance quality and strong genre controllability. It consists of two stages: (1) High-Fidelity Dance Quantization Stage (HFDQ), which encodes dance motions into a latent representation by Finite Scalar Quantization (FSQ) and reconstructs them with kinematic-dynamic constraints, and (2) Genre-Aware Dance Generation Stage (GADG), which maps music into the latent representation by synergistic utilization of Mixture-of-Experts (MoE) mechanism with Mamba-Transformer hybrid backbone. Extensive experiments on the FineDance and AIST++ dataset demonstrate the state-of-the-art performance of MEGADance both qualitatively and quantitatively. Code will be released upon acceptance.
1 Introduction
Music-to-dance generation is a crucial task that translates auditory input into dynamic motion, with significant applications in virtual reality, choreography, and digital entertainmentli2021ai ; li2023finedance . By automating this process, it enables deeper exploration of the intrinsic relationship between music and movementyang2024beatdance , while expanding possibilities for creative content generation. Due to its broad impact, music-to-dance generation has attracted increasing attentionsiyao2022bailando ; siyao2023bailando++ ; li2024lodge++ .
Current music-to-dance generation approaches have witnessed rapid progress and can be broadly categorized into two paradigmssiyao2022bailando ; zhuang2023gtn : (1) One-stage methods directly map musical features to human motionli2024lodge ; li2024lodge++ ; huang2020dance . (2) Two-stage methods first construct choreographic units and then learn their probability distributions conditioned on music siyao2022bailando ; siyao2023bailando++ ; siyao2024duolando . However, previous methods only treat genre as a weak auxiliary bias rather than the core semantic driverli2024exploring ; li2023finedance ; li2024lodge , facing several essential problems such as misaligned music-motion synchronization and disrupted dance genre continuity. For example, A Uyghur movements clip is inappropriately mixed into the Popping routine, when a typical Popping music exhibits complex transitions in rhythm and intensity, as shown in Fig. 1. This oversight leads to unsatisfactory visual effects, and fails to meet genre-specific user demands.
Figure 1: MEGADance enhances choreography consistency by decoupling it into dance generality and genre specificity via the Mixture-of-Experts design. Compared to previous methods, it produces synchronized dance with genre continuity, even under complex music conditions.
To address these limitations, we propose MEGADance, the first Mixture-of-Expertsshazeer2017outrageously (MoE) architecture for Genre-Aware 3D Dance Generation. By decoupling choreographic consistency into dance generality, modeled by Universal Experts shared across all genres, and genre specificity, captured by Specialized Experts selected via genre-guided hard routing, MEGADance enables robust musical alignment and fine-grained stylistic fidelity. Due to its structured inductive bias, MEGADance exhibits strong genre control and remains robust even in the presence of modality conflicts, such as generating a Breaking dance for soft-paced Chinese music while preserving rhythmic and dynamic alignment.
Specifically, MEGADance comprises two stages. (1) High-Fidelity Dance Quantization Stage (HFDQ), which encodes dance motions into a latent representation. In the HFDQ stage, we introduce Finite Scalar Quantizationmentzer2023finite (FSQ), which replaces traditional VQ-VAE codebooks to mitigate codebook collapse and enhance latent diversity. Additionally, we impose kinematic constraints by simultaneously reconstructing 3D joints through Forward Kinematicsloper2023smpl from SMPLloper2023smpl , and dynamic constraints by reconstructing dances while considering velocity and acceleration, to enhance spatio-temporal coherence. (2) Genre-Aware Dance Generation Stage (GADG), which maps music into the latent representation. In the GADG, the Universal Experts and Specialized Experts work synergistically to jointly modeling dance generality and genre specificity. For genre-disentangled expert design, Universal Experts model universal rhythmic and temporal structures across all genres, enhancing robustness to diverse musical inputs and stabilizing cross-modal alignment; and Specialized Experts specialize in fine-grained stylistic variations unique to each genre, guided by hard routing to effectively disentangle genre-dependent features from genre-invariant dynamics. For expert structure, each expert, whether Specialized or Universal, adopts an autoregressive Mamba-Transformer hybrid backbone, combining Mamba’s efficient intra-modal local dependency capture gu2023mamba with the Transformer’s cross-modal global contextual understandingvaswani2017attention , thereby enabling the generation of temporally coherent and musically aligned dance motions.
The contributions of our work can be summarized as: (1) We introduce MEGADance, the first Mixture-of-Experts (MoE) architecture for music-to-dance generation, designed to enhance choreographic consistency by decoupling it into dance generality and genre specificity. MEGADance achieves state-of-the-art (SOTA) performance and demonstrates robust genre controllability, as demonstrated through extensive qualitative and quantitative experiments on the AIST++li2021ai and FineDanceli2023finedance datasets. (2) We propose a High-Fidelity Dance Quantization framework that introduces FSQ with kinematic-dynamic dual constraints, ensuring complete codebook utilization (100% vs. VQ-VAEoord2017neural ’s 75%) while achieving excellent reconstruction accuracy. (3) We design a Mamba-Transformer hybrid backbone for music-to-dance generation, combining Mamba’s efficient intra-modal local dependency capture with the Transformer’s cross-modal global contextual understanding.
2 Related Work
2.1 One-Stage Music-to-Dance Generation
Music and dance are deeply interconnected, leading to significant advancements in the field of music-driven 3D dance generation. Researchers utilize musical features extracted via tools like Librosali2021ai , Jukeboxtseng2023edge , and MERTyang2024codancers to predict human motion, including SMPLloper2023smpl parametersli2021ai and body keypointssiyao2022bailando . Early methods primarily employ encoder-decoder architectures to directly obtain entire human motion sequence lee2018listen ; tang2018dance ; le2023music ; correia2024music ; li2021ai . Recognizing the natural hierarchical structure of human joints, some researchers introduced Graph Convolutional Networks (GCNs)ferreira2021learning ; yan2019convolutional to enhance interaction at the joint level, thereby improving the biomedical plausibility of the generated motions. In AIGC, Generative Adversarial Networks (GANs) are widely applied, some researchers introduced it in music-to-dance taskschen2021choreomaster ; huang2020dance ; yang2024cohedancers . Specifically, GANs’ generators produce dance motions from music, with discriminators providing feedback to guide generated motions more natural. Recently, Diffusion Models have shown remarkable success in various AIGC tasks, with notable applications extending to the music-to-dance domaintseng2023edge ; li2023finedance ; li2024lodge ; li2024lodge++ ; le2023controllable , but the computational cost of the sampling process remains high, especially in long-sequence generation scenarios for the music-to-dance task. However, the lack of explicit constraints to maintain the generated pose within proper spatial boundaries often leads to nonstandard poses that extend beyond the dancing subspace during inference, resulting in low dance quality in practical applications.
2.2 Two-Stage Music-to-Dance Generation
Leveraging the inherent periodicity of dance kinematics, researchers propose Two-Stage methods, including (1) Dance Quantization stage: curating choreographic units from motion databases, and (2) Dance Generation stage: learning music-conditioned probability distributions over these units. As these choreographic units are derived from real human motion data, two-stage approaches naturally benefit from a biomechanical plausibility prior, contributing to the realism of generated dances.
Dance Quantization Stage. Traditional methods ye2020choreonet ; aristidou2022rhythm ; huang2022genre ; chen2021choreomaster construct choreographic units through uniform segmentation of motion sequence, incurring high computational overhead. Recent works gong2023tm2d ; guo2022generating employ VQ-VAE for intelligent unit construction, significantly reducing time/space complexity. Considering the relative independence of upper and lower body movements, siyao2022bailando ; siyao2023bailando++ construct choreographic unit for lower and upper parts separately, improving motion reconstruction through expanded unit capacity ($L\to L\times L$). However, above works predominantly operate on 3D human body keypoints, which lack expressiveness in capturing nuanced motion details. siyao2023bailando++ ; li2024lodge++ construct choreographic units in the SMPL pose space but apply uniform treatment across joints, neglecting the body’s kinematic hierarchy, such as root errors propagate globally through kinematic chains while hand errors remain localized.
Dance Generation Stage. To model choreographic unit distributions, Choreomaster chen2021choreomaster employs a GRU-based backbone, while DanceRevolution huang2020dance uses RNNs. Recent works like Bailando siyao2022bailando and Bailando++siyao2023bailando++ adopt cross-modal Transformers for improved temporal modeling and music-motion alignment. Moreover, Everything2Motion fan2024everything2motion and TM2D gong2023tm2d leverage pretrained models in text-to-motionguo2022generating ; guo2022tm2t generation to improve motion quality, often at the expense of choreographic complexity and creativity. To enrich input representations, li2023finedance ; li2024lodge ; huang2022genre ; zhuang2023gtn introduce genre information via shallow fusion, such as cross attention huang2022genre or feature addition zhuang2023gtn . However, these approaches remain insufficient for achieving robust genre controllability, particularly under cross-modal conflicts, such as generating Breaking dances conditioned on slow-tempo traditional Chinese music.
In conclusion, existing methods face two key limitations: (1) VQ-VAE-based quantization suffers from low codebook utilization; (2) Insufficient utilization of genre information results in poor music-motion synchronization and disrupted dance genre continuity. Thus, we propose MEGADance, a two-stage framework. The Dance Quantization stage employs FSQ with kinematic and dynamic constraints to enhance codebook efficiency while preserving reconstruction fidelity. The Dance Generation stage adopts an MoE architecture with a Mamba-Transformer backbone to jointly capture dance generality and genre specificity with efficiency.
Figure 2: Overview of MEGADance. MEGADance employs FSQs with kinematic-dynamic constraints for body-part reconstruction in HFDQ, coupled with a MoE-based Mamba-Transformer architecture that generates music-aligned latent representations in GADG.
3 Methodology
3.1 Problem Definition
Given a music sequence $M=\{m_{0},m_{1},...,m_{T}\}$ and a dance genre label $g$, our objective is to synthesize the corresponding dance sequence $S=\{s_{0},s_{1},...,s_{T}\}$, where $m_{t}$ and $s_{t}$ denotes the music and dance feature at time step $t$. We define each music feature $m_{t}$ as a 35-dim vectorli2024lodge extracted by Librosamcfee2015librosa , including 20-dim MFCC, 12-dim Chroma, 1-dim Peak, 1-dim Beat and 1-dim Envelope. We encode genre label $g$ as a one-hot vector. We represent each dance feature as a 147-dim vector $s_{t}=[\tau;\theta]$, where $\tau$ and $\theta$ encapsulate the root translation and 6-dim rotation representationzhou2019continuity of the SMPLloper2023smpl model, respectively. Furthermore, we synchronize the music sequence with the dance sequence at a temporal granularity of 30 frames per second.
3.2 High-Fidelity Dance Quantization
Finite Scalar Quantization with Motion Decomposition. Choreographic units serve as the fundamental building blocks of dance composition, forming the basis for structuring and connecting movements. Despite variations in style and tempo, dances across genres exhibit common underlying units. Our objective is to unsupervisedly encapsulate these units into a versatile and reusable codebook, enabling any dance sequence to be represented as a sequence of discrete codebook elements. To account for the relative independence between upper-body and lower-body movements during dance, we maintain separate codebooks for the upper and lower $\mathcal{Z}=\{\mathcal{Z}^{u}_{k},\mathcal{Z}^{l}_{k}\}$ body, where $k$ represent codebook size. Additionally, root translation velocities are associated with the lower body to preserve natural motion dynamics. This decomposition allows the combination of different code pairs to cover a wider array of choreographic units.
Our 3D motion reconstruction approach, illustrated in Fig. 2, initiates with a Dance Encoder $\mathbf{E}$ (a three-layer 1D-CNN for information aggregation and a two-layer MLP for dimension adjustment) encoding the dance sequence $S=\{S^{u},S^{l}\}$ into context-aware features $\mathbf{z}=\{\mathbf{z^{u}},\mathbf{z^{l}}\}$. These features are quantized using Finite Scalar Quantization (FSQ) to obtain $\hat{\mathbf{z}}=\{\hat{\mathbf{z^{u}}},\hat{\mathbf{z^{l}}}\}$, which are then decoded by Dance Decoder $\mathbf{D}$ (a two-layer MLP for dimension adjustment and a three-layer 1D TransConv for information restoration) to reconstruct the dance movement $\hat{S}=\{\hat{S^{u}},\hat{S^{l}}\}$. To resolve the codebook collapse problem caused by the conventional VQ-VAEvan2017neural based quantization, we adopt FSQ. By replacing the discrete "argmin" codebook selection with scalar quantization via differentiable bounded rounding, FSQ enables balanced utilization and stable gradient propagation:
$\hat{\mathbf{z}}=f(\mathbf{z})+\text{sg}\left[\text{Round}[f(\mathbf{z})]-f(\mathbf{z})\right],$ (1)
where $f(\cdot)$ is the bounding function, setting as the $\mathrm{sigmoid}(\cdot)$ function in our practice. Each channel in $\hat{\mathbf{z}}$ will be quantized into one of the unique $L$ integers, therefore we have $\hat{\mathbf{z}}\in\{1,\dots,L\}^{d}$. The codebook size is calculated as $k=\prod_{i=1}^{d}L_{i}$, and $L,d$ are super parameter. In conclusion, FSQ with Motion Decomposition expands its effective motion representation capacity, thereby enhancing the diversity of subsequent generated dance.
Motion Reconstruction with Kinetic-Dynamic Constraint. Unlike VQ-VAE requiring additional loss to update any extra lookup codebook, FSQ directly integrates numerical approximations "round" within its workflow. The Dance encoder $E$ and decoder $D$ are simultaneously learned with the codebook via the following loss function:
$\mathcal{L}_{FSQ}=\mathcal{L}_{\text{smpl}}(\hat{S},S)+\mathcal{L}_{\text{joint}}(\hat{J},J).$ (2)
The $\mathcal{L}_{\text{smpl}}$ is the reconstruction loss that ensures the predicted 3D SMPL sequence closely aligns with the ground truth. Simple reconstruction on SMPL parameters treats all joints equally, neglecting the complex hierarchical tree structure of human body joints, different joints vary in their tolerance to errors. For instance, errors at the root node propagate throughout all nodes, whereas errors at the hand node primarily affect only itself. Thus, we execute Forward Kineticloper2023smpl techniques to derive 3D joints and apply reconstruction constraints $\mathcal{L}_{\text{joint}}$ between $\hat{J}$ and $J$. Moreover, the reconstruction loss accounts not only for the spatial positions but also for the velocities ($\alpha_{1}$) and accelerations ($\alpha_{2}$) of the movements:
$\mathcal{L}_{\text{smpl}}(\hat{S},S)=\|\hat{S}-S\|_{1}+\alpha_{1}\|\hat{S}^{\prime}-S^{\prime}\|_{1}+\alpha_{2}\|\hat{S}^{\prime\prime}-S^{\prime\prime}\|_{1},$ (3)
$\mathcal{L}_{\text{joint}}(\hat{J},J)=\|\hat{J}-J\|_{1}+\alpha_{1}\|\hat{J}^{\prime}-J^{\prime}\|_{1}+\alpha_{2}\|\hat{J}^{\prime\prime}-J^{\prime\prime}\|_{1}.$
Through training, our method facilitates the interchangeability of orthographic memory codes, enabling the synthesis of new motions from existing choreographic units by recombining different code elements.
3.3 Genre-Aware Dance Generation
With dance sequences represented as discrete latent codes, the music-to-dance generation task is simplified from a regression problem into a classification problem, where the goal is to select appropriate pose codes from a codebook rather than predict continuous motion parameters.
3.3.1 Mixture-of-Experts Architecture
As illustrated in Fig. 2, we perform cross-modal autoregressive generation. Given music features $m_{1:T}$ extracted using Librosa, dance genre label $g$, and the previous pose codes $p_{0:T-1}=\{p_{0:T-1}^{l},p_{0:T-1}^{u}\}$ encoded by the Motion Tokenization $\mathbf{E}$ in HFDQ, the GADG predicts action probabilities $a_{0:T-1}=\{a_{0:T-1}^{l},a_{0:T-1}^{u}\}$ of every $z_{i}\in\mathcal{Z}$, using an $L$-layer MoE architecture. To align the predicted action probabilities $a_{0:T-1}$ with the next pose codes $p_{1:T}$, we employ a supervised Cross-Entropysiyao2022bailando loss, where each predicted action $a_{t}$ is matched to its corresponding target pose code $p_{t+1}$. The inference of GADG includes: 1) Short sequences ($\leq 5.5s$) via autoregressive generation, 2) Long sequences via sliding-window prediction with 5.5s overlap.
Specifically, each MoE layer contains a Specialized Expert and a Universal Expert, which jointly model dance generality and genre specificity. Specialized Experts (e.g., Pop Expert, Jazz Expert) are conditionally activated based on the genre label $g$, and input features are routed to the corresponding expert via a hard routing mechanism. In parallel, features from all genres are processed by the shared Universal Expert to capture genre-invariant dynamics. For expert structure, each expert, whether Specialized or Universal, adopts an autoregressive Mamba-Transformer hybrid backbone: Mamba captures intra-modal local dependencies, while Transformer encodes cross-modal global context, thereby enabling the generation of temporally coherent and musically aligned dance motions.
3.3.2 Genre-Disentangled Experts Design
Specialized Experts. The Specialized Experts are designed to capture genre-specific stylistic patterns, motivated by two core considerations: (1) Structural Inductive Bias: By isolating parameters across experts, the model enforces separation of genre-specific motion motifs (e.g., Krump’s grounded explosiveness vs. Contemporary’s fluid transitions), thereby preserving distinct stylistic representations. This separation also introduces genre-aware control priors that mitigate cross-genre interference, which is critical for genre-conditioned dance generation. (2) Computational Efficiency: Leveraging sparse MoE design Shazeer2017 , each input is routed to a single expert, significantly reducing parameter redundancy and computational cost.
Universal Experts. The Universal Expert learns generalizable representations to complement Specialized Experts through two key roles: (1) Fundamental Choreographic Prior: It learns shared low-level patterns across genres (e.g. periodicity, beat synchronization, and biomechanical consistency). In contrast, models relying solely on Specialized Experts often fail under modality mismatch (e.g., producing static or repetitive movements when Ballet music is processed by a Popping Expert). The Universal Expert provides a genre-agnostic prior that enhances stability and expressiveness under complex input conditions. (2) Disentangled Representation: By disentangling shared and genre-specific factors, the model allows each expert to specialize in distinct subspaces, enhancing generation qualitylocatello2020weakly .
3.3.3 Mamba-Transformer Hybrid Backbone
Cross-Modal Global-Context Modeling. Leveraging the Transformer’s global receptive field vaswani2017attention , we concatenate cross-modal features along the temporal axis and facilitate structured interactions among music, upper-body, and lower-body representations, by a Attention layer and a Feed Forward layer. The attention layer is the core component that defines the computational dependencies among sequential data elements and is implemented as:
$\text{Attention}(Q,K,V,M)=\text{softmax}\left(\frac{QK^{T}+M}{\sqrt{C}}\right)V,$ (4)
where $Q,K,V$ denote the query, key, and value from input, and $M$ is the mask, which determines the type of attention layers. The two most common attention types are full attention vaswani2017attention , which enables global context exchange, and causal attention radford2018improving , which restricts each position to attend only to current and past inputs. Given that music-to-dance generation is typically applied to long sequence while being constrained by limited computational resources, training is commonly conducted on short clips. During inference, the sequence is first extended autoregressively up to the training length (step 1), and then completed using a sliding window approach for the remaining part (step 2). Training driven by standard causal attention only aligns with step 1 and fails to account for the dominant step 2 during inference, thereby limiting generation performance. To better align training with inference, we introduce a sliding-window attention mechanism that mimics the generation process. The attention mask $M\in\mathbb{R}^{3T\times 3T}$ is structured as a $3\times 3$ block matrix, where each block is a $T\times T$ sliding-window mask, enabling cross-modal global-context attention.
Intra-Modal Local-Dependency Modeling. While the Transformer excels at temporal modeling, it is inherently position-invariant and captures sequence order only through positional encodings vaswani2017attention , which limits its deep understanding of local dependencies. In contrast, music-to-dance generation demands strong local continuity between movements. Owing to its inherent sequential inductive bias, Mamba gu2023mamba has demonstrated strong performance in modeling fine-grained local dependencies xu2024mambatalk ; fu2024mambagesture . We therefore apply independent Mamba to the music, upper-body, and lower-body features respectively, to model their intra-modal local dependencies. Specifically, Selective State Space model (Mamba) incorporates a Selection mechanism and a Scan module (S6) gu2023mamba to dynamically select salient input segments for efficient sequence modeling. Unlike the traditional S4 model gu2021efficiently with fixed parameters $A,B,C$, and scalar $\Delta$, Mamba adaptively learns these parameters via fully-connected layers, enhancing generalization capabilities. Mamba employs structured state-space matrices, imposing constraints that improve computational efficiency. For each time step $t$, the input $x_{t}$, hidden state $h_{t}$, and output $y_{t}$ follow:
$h_{t}$ $=\bar{A}_{t}h_{t-1}+\bar{B}_{t}x_{t},$ (5)
$y_{t}$ $=C_{t}h_{t},$
where $\bar{A}_{t},\bar{B}_{t},C_{t}$ are dynamically updated parameters. Through discretization with sampling interval $\Delta$, the state transitions become:
$\bar{A}$ $=\exp(\Delta A),$ (6)
$\bar{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B,$
$h_{t}$ $=\bar{A}h_{t-1}+\bar{B}x_{t},$
where $(\Delta A)^{-1}$ is the inverse of $\Delta A$, and $I$ denotes the identity matrix. The scan module captures temporal dependencies by applying trainable parameters across input segments.
4 Experiment
4.1 Dataset
FineDance. FineDance li2023finedance is the largest public dataset for 3D music-to-dance generation, featuring professionally performed dances captured via optical motion capture. It provides 7.7 hours of motion data at 30 fps across 16 distinct dance genres. Following li2024lodge , we evaluate on 20 test-set music clips, generating 1024-frame (34.13s) dance sequences.
AIST++. AIST++ li2021ai is a widely used benchmark comprising 5.2 hours of 60 fps street dance motion, covering 10 dance genres. Following li2021ai , we use 40 test-set music clips to generate 1200-frame (20.00s) sequences.
4.2 Quantitative Evaluation
Table 1: Comparison with SOTAs on the FineDance dataset.
Quality Creativity Alignment User Study
[…]
GT 0 0 0 10.98 7.45 6.07 0.215 4.39 4.35 4.48
Bailando++siyao2023bailando++ 54.79 16.29 8.42 6.18 5.98 4.73 0.213 3.85 3.50 3.82
FineNetli2023finedance 65.15 23.81 13.22 5.84 5.19 4.29 0.219 3.62 3.65 3.47
Lodgeli2024lodge 55.03 14.87 5.22 6.14 6.18 5.50 0.218 4.18 4.17 4.08
MEGADance 50.00 13.02 2.52 6.23 6.27 5.78 0.226 4.25 4.30 4.23
Comparison. We evaluate MEGADance against state-of-the-art (SOTA) baselines on both the FineDance and AIST++ datasets using a comprehensive suite of metrics. For each generated sequence, we compute Fréchet Inception Distance ($FID$) and Diversity ($DIV$) across three feature spaces: (1) Kinetic (k), capturing motion dynamics; (2) Geometric (g), encoding spatial joint relations; and (3) Style (s), extracted via a Transformer-based genre classifier. We also assess music-motion synchronization using Beat Align Score ($BAS$), following li2021ai ; li2024lodge . On the FineDance dataset (Tab. 1), MEGADance outperforms all baselines, achieving the lowest FID in all three feature types ($FID_{k}$=50.00, $FID_{g}$=13.02, $FID_{s}$=2.52), the highest motion diversity ($DIV_{k}$=6.23, $DIV_{g}$=6.27, $DIV_{s}$=5.78), and the best BAS (0.226). On the AIST++ datasets (Tab. 4.3), MEGADance again ranks first in FID ($FID_{k}$=25.89, $FID_{g}$=12.62), and achieves strong performance in diversity ($DIV_{g}$=5.84, $DIV_{s}$=6.23) and BAS (0.238). These results underscore the effectiveness of our genre-aware Mixture-of-Experts design in balancing motion quality, creativity, and synchronization across diverse datasets.
User Study. Dance’s inherent subjectivity makes user feedback essential for evaluating generated movementslegrand2009perceiving in the music-to-dance generation task. We select 30 in-the-wild music segments (34 seconds each) and generate dance sequence using above models. These sequences are evaluated through a double-blind questionnaire, by 30 participants with backgrounds in dance practice, including undergraduate and graduate-level students. The questionnaires are based on a 5-point scale (Great, Good, Fair, Bad, Terrible) and assess three aspects: Dance Synchronization ($DS$, alignment with rhythm and style), Dance Quality ($DQ$, biomechanical plausibility and aesthetics), and Dance Creativity ($DC$, originality and range). As shown in Tab. 1, MEGADance significantly outperforms all baselines across user-rated metrics (i.e., $DS=4.30$, $DQ=4.25$, $DC=4.23$). Its high scores in various aspects underscore its superiority in generating movements in terms of human preferences.
Performance Efficiency. In addition to accuracy, we evaluate the runtime efficiency of MEGADance. The model generates one second of feedback in just 0.19 seconds, highlighting its suitability for real-time applications. This speed enables seamless integration into interactive systems, where rapid feedback is crucial for user engagement in practice.
4.3 Qualitative Evaluation
To assess the visual quality of the generated dance sequences, we perform a qualitative comparison between MEGADance and several existing baseline models, as depicted in Fig. 3. In terms of expressiveness, MEGADance outperforms the competing methods in several key areas. For instance, Lodgeli2024lodge struggles with stylistic consistency, often blending conflicting dance genres, such as incorporating Uyghur movements into a typical Breaking routine, leading to a disjointed aesthetic. FineNetli2023finedance , while capable of generating movement sequences, suffers from significant artifacts, including unnatural sliding and teleportation, which detract from the fluidity and physical realism of the motions. Additionally, Bailando++siyao2023bailando++ demonstrates a lack of diversity, with movements frequently repeating and a heavy reliance on static, in-place gestures, limiting the range of expressive movement patterns. These findings underscore the superiority of MEGADance in generating diverse, genre-consistent, and musically synchronized dance sequences.
Figure 3: Qualitative Analysis on a typical Breaking Battle music clip.
Figure 4: Visualization of Genre Controllability on a representative Chinese music clip.
Table 2: Comparison on the AIST++ Dataset.
[…]
GT 0 0 9.04 7.52 0.232
FACTli2021ai 35.35 22.11 5.94 6.18 0.221
Bailando++siyao2023bailando++ 30.21 15.48 5.35 5.13 0.228
EDGEtseng2023edge 42.16 22.12 3.96 4.61 0.233
Lodgeli2024lodge 35.72 17.92 5.72 5.91 0.247
MEGADance 25.89 12.62 5.84 6.23 0.238
Table 3: Comparison for Genre Controllability.
[…]
GT 0 6.07 78.31 76.35
FineNetli2023finedance 13.22 4.29 42.06 37.44
Lodgeli2024lodge 5.22 5.50 51.86 45.23
MEGADance 2.52 5.78 75.64 70.81
4.4 Genre Controllability Evaluation
To quantitatively assess genre controllability, we compare MEGADance with Lodge li2024lodge and FineNet li2023finedance . Using ground-truth genre labels, we evaluate style alignment ($FID_{s}$) and style diversity ($DIV_{s}$) on 20 test clips. We further assess genre classification accuracy ($ACC$) and F1 score ($F1$), conditioned on the correct genre and four randomly sampled negative genres. As shown in Tab. 4.3, MEGADance achieves the best performance across all metrics. It significantly reduces $FID_{s}$ (2.52) while improving diversity ($DIV_{s}$ = 5.78). Despite potential cross-modal conflicts (e.g., assigning a Popping genre to a typical Chinese music clip), MEGADance achieves high genre discriminability ($ACC$ = 75.64, $F1$ = 70.81), closely approaching ground-truth performance. Compared to FineNet and Lodge, our method produces motion sequences that are both more stylistically coherent and genre-distinctive. Our MoE-based genre routing prevents cross-genre interference via disentangled expert subspaces activated by discrete labels, whereas naive continuous fusion (e.g. Feature Addition in li2023finedance ; li2024lodge or Cross Attention in huang2022genre ) inherently blurs stylistic boundaries.
To explore genre controllability from a visual perspective, we assign different dance genres (distal: Breaking/Locking, proximal: Dunhuang/Uyghur) to a representative Chinese music clip. We recommend watching the supplementary video for more details. As shown in Fig. 4, the generated motions exhibit both genre fidelity and music synchrony: (1) Breaking: agile footwork with rapid steps and directional shifts, driven by percussive rhythms and dynamic weight transfers; (2) Locking: exaggerated arm swings and torso isolations, punctuated by syncopated, guitar-mimicking gestures; (3) Dunhuang: fluid upper-body arcs with slow rotations and knee undulations, mirroring melodic phrasing and visual symmetry; (4) Uyghur: rapid spins with hand-to-face motifs, emphasizing rotational clarity and rhythmic precision.
Table 4: Ablation study of the two-stage MEGADance architecture on the FineDance dataset.
[…]
GT 0 0 0 0.215
w/o SE 53.05 19.26 7.95 0.218
w/o UE 54.50 15.52 2.91 0.223
w/o Mamba 56.29 14.51 2.67 0.221
Ours 50.00 13.02 2.52 0.226
(a) High-Fidelity Dance Quantization Stage.
SMPL Joint
[…]
w/o Kin. Loss 0.0238 0.0847 0.0089 0.0507
w/o Dyn. Loss 0.0201 0.0779 0.0073 0.0482
FSQ $\rightarrow$ VQ-VAE 0.0308 0.0984 0.0220 0.0842
Ours 0.0200 0.0770 0.0069 0.0469
(b) Genre-Aware Dance Generation Stage.
4.5 Ablation Study
4.5.1 Genre-Aware Dance Generation Stage
We conduct an ablation study to evaluate the contribution of three core components in the Genre-Aware Dance Generation stage: Specialized Experts (SE), Universal Experts (UE), and the Intra-Modal Local-Dependency Modeling (Mamba), with results summarized in Tab. 4(a). We recommend watching the supplementary video. (1) Specialized Experts (SE). Replacing the SE results in substantial performance degradation across all metrics, especially on $FID_{s}$ (7.95 vs. 2.52), confirming its critical role in preserving stylistic fidelity across genres. (2) Universal Expert (UE). Removing the UE leads to clear drops in $FID_{k}$ (54.50 → 50.00) and $FID_{g}$ (15.52 → 13.02), while having only minor impact on $FID_{s}$ and $BAS$. This suggests the UE’s effectiveness in providing generalizable priors that enhance structural and dynamic consistency. (3) Mamba. Replacing the Mamba in backbone results in moderate performance declines on all metrics (e.g., $FID_{k}$ from 50.00 to 56.29), demonstrating Mamba’s advantage in modeling fine-grained local dependencies and improving overall motion quality.
4.5.2 High-Fidelity Dance Quantization Stage
We investigate the effectiveness of three key components in High-Fidelity Dance Quantization Stage: Finite Scalar Quantization (FSQ), the Kinematic Loss ($\mathcal{L}_{\text{Kin.}}$), and the Dynamic Loss ($\mathcal{L}_{\text{Dyn.}}$). Tab. 4(b) reports the $MSE$ and $MAE$ on both SMPL parameters and 3D joint positions. (1) FSQ. Replacing FSQ with VQ-VAE leads to a significant performance drop in all metrics (e.g., Joint $MSE$ increases from 0.0069 to 0.0220), validating FSQ’s superiority. Moreover, replacing VQ-VAE with FSQ achieves full codebook utilization (75% → 100%). (2) Kinematic Loss. Removing the kinematic loss notably increases errors in both SMPL ($MSE$: $0$ 0.0238) and joint space ($MAE$: 0.0469 → 0.0507), highlighting its role in enforcing accurate structural constraints via forward kinematics. (3) Dynamic Loss. Excluding the dynamic loss results in a moderate degradation in temporal fidelity (e.g., Joint $MSE$: 0.0069 → 0.0073), demonstrating its contribution for temporal fidelity.
5 Conclusion
In this paper, we present MEGADance, a genre-aware MoE-based architecture for music-to-dance generation. MEGADance enhances choreography consistency by decoupling it into dance generality and genre specificity via an MoE design. Through the synergy of high-fidelity dance quantization stage and genre-adaptive dance generation stage, MEGADance achieves state-of-the-art performance and strong genre controllability. In future work, we plan to extend MEGADance with text conditioning to enable more interactive and flexible dance generation.
Appendix
A.1 Implementation Details
High-Fidelity Dance Quantization. In the High-Fidelity Dance Quantization Stage, we use a shared codebook configuration for the upper and lower body branches. The model is trained on 8-second SMPL 6D rotation sequences sampled at 30fps, where $S,\hat{S}\in\mathbb{R}^{T\times 147}$ (i.e., $T=240$). For data construction, we augment the training set using a sliding window approach with a window size of 240 and a stride of 16. A three-layer CNN encoder $E$ performs temporal downsampling, and a three-layer transposed convolution decoder $D$ performs upsampling. The latent codes for the lower and upper body are $p^{l},p^{u}\in\mathbb{R}^{T^{\prime}}$, with $T^{\prime}=30$. In the Finite Scalar Quantization module, the codebook size is 4375, with $L=[7,5,5,5,5]$, and the feature dimension is set to 512. For reconstruction, we use both SMPL-parameter loss $\mathcal{L}_{\text{smpl}}$ and joint-position loss $\mathcal{L}_{\text{joint}}$, with velocity and acceleration terms weighted by $\alpha_{1}=0.5$ and $\alpha_{2}=0.25$, respectively. The model is trained for 200 epochs using the Adam optimizer, with exponential decay rates of 0.5 and 0.99 for the first and second moment estimates. A fixed learning rate is used with a batch size of 32. The experimental setup is consistent across FineDance and AIST++.
Genre-Aware Dance Generation. In the Genre-Aware Dance Generation Stage, we adopt a Mamba-Transformer hybrid architecture, trained on latent codes $p^{l},p^{u}\in\mathbb{R}^{30}$ extracted from the High-Fidelity Dance Quantization Stage, using 8-second dance sequences at 30fps. For data construction, we augment the training set using a sliding window approach with a window size of 240 and a stride of 16. In MEGADance, the Music Encoder consists of $L=6$ processing layers. The Mamba block is configured with a model dimension of 512, state size of 16, convolution kernel size of 4, and expansion factor of 2. The Transformer block uses a hidden size of 512, 8 attention heads, a feedforward dimension of 2048, and a dropout rate of 0.25. For Slide Window Attention, we set the autoregressive step to 22 and the sliding window step to 8 to construct the attention matrix. For input representation, genre labels (16 classes from FineDance) are embedded using nn.Embedding to match the 512-dimensional latent space, while music features extracted by Librosa (35 dimensions) are projected to 512 dimensions via a two-layer MLP. For output, MEGADance predicts 4375-class distributions via softmax for upper-body and lower-body codebook respectively. The model is optimized using Adam with exponential decay rates of 0.9 and 0.99 for the first and second moment estimates, respectively, trained for 80 epochs with a fixed learning rate and a batch size of 64. The experimental setup is consistent across FineDance and AIST++.
A.2 Qualitative Analysis for Ablation Study
Figure 5: Qualitative Analysis for Ablation Study. MEGADance generates visually expressive dance motions, outperforming others in terms of stylistic consistency and movement diversity.
In this section, we conduct a qualitative analysis to evaluate the contribution of each component in the Genre-Aware Dance Generation stage. As illustrated in Fig. 5, the Specialized Experts (SE), Universal Experts (UE), and the Mamba-enhanced backbone (Mamba) each play a crucial role in shaping the quality of the generated dance motions.
Removing SE, which is responsible for capturing genre-specific stylistic features, results in a significant loss of genre identity. This removal leads to mismatches where, for example, soft or fluid movements are applied to intense, percussive music, breaking the stylistic coherence expected for the genre. In contrast, the exclusion of UE impacts the overall complexity of the generated motion. Without UE, the generated sequences tend to be overly simplistic, often consisting of static poses or repetitive, monotonous movements, such as constant hand-raising, which lack the dynamic variation essential for engaging dance sequences. Furthermore, omitting the Mamba module, which enhances the backbone with a selection mechanism and scan module, results in a significant decrease in both movement diversity and alignment with the music. The generated dances become less responsive to the rhythmic and dynamic changes in the music, leading to sequences that feel disjointed or fail to reflect the musical structure accurately.
Collectively, these observations highlight the importance of each component in the overall framework. The combination of SE, UE, and Mamba ensures that the generated dance is not only genre-appropriate but also rich in motion variety and tightly aligned with the music.
Figure 6: The screenshots of user study website for participants.
A.3 Questionnaire for User Study
User feedback is essential for evaluating generated dance movements in the music-to-dance generation task, due to the inherent subjectivity of dance [45]. We select 30 real-world music segments, each lasting 34 seconds, and generated dance sequences using the models described above. These sequences are evaluated through a double-blind questionnaire completed by 30 participants with dance backgrounds, including undergraduate and graduate students. Participants are compensated at a rate exceeding the local average hourly wage. The questionnaires used a 5-point scale (Great, Good, Fair, Bad, Terrible) to assess three aspects: Dance Synchronization (DS, alignment with rhythm and style), Dance Quality (DQ, biomechanical plausibility and aesthetics), and Dance Creativity (DC, originality and range).
The screenshot of our user study website is shown in Fig. 6, displaying the template layout presented to the participants. In addition to the main trials, participants are also subjected to several catch trials, which involved displaying Ground Truth videos and videos with distorted motion. Participants who failed to rate the GT videos higher and the distorted motion videos lower are considered unresponsive or inattentive, and their data were excluded from the final evaluation.
A.4 Future Work
Customized Dance Generation While our current work successfully enables genre-aware control in music-to-dance generation, genre labels inherently impose rigid constraints and offer limited flexibility for user intent expression. Existing controllable generation approaches remain insufficiently expressive for practical deployment [31, 11, 7]. In future work, we plan to extend control modalities beyond predefined genre categories by incorporating free-form textual descriptions. Compared to genre labels, text allows users to articulate choreography requirements in a more intuitive and nuanced manner, facilitating personalized and expressive dance generation. This direction not only enhances user interactivity and creativity but also opens up new opportunities for content-driven applications in virtual performance and human-computer interaction.
Noise-Resistant Dance Generation 3D motion capture data often suffer from noise artifacts such as sudden positional jumps or temporal discontinuities, as observed even in high-quality datasets like FineDance[2]. Moreover, the limited scale of 3D dance datasets makes models prone to overfitting. Future research should explore robust architectures and data augmentation strategies that maintain motion plausibility and stylistic coherence under noisy or incomplete input, thereby improving the reliability and generalization of music-to-dance generation systems.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Overall, the abstract and introduction provide a concise yet comprehensive summary of the paper’s objectives, methods, and findings, accurately reflecting its contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification:Yes, we discuss the limitations in section App. 4.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: By adhering to the principles mentioned in the Guidelines, we ensures that each theoretical result is underpinned by a full set of assumptions and complete, correct proofs, thus reinforcing the credibility and reliability of the paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: In section Sec. 4 and App.1, we report all the experiments setting, implementation details and metrics, which disclose all the information needed to reproduce the main experimental results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code is available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In section Sec. 4 and App. 1, we report all the experiments setting and implementation details, facilitating readers’ understanding of the results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: While extensive quantitative and user study results are provided in Sec. 4, the paper does not report error bars, variance, or statistical tests for the quantitative metrics or user study scores.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We show model runtimes in detail in Section 4.1 Efficiency Analysis.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we do.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The work we perform makes no society impact. It is only an academic study.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly cite the original owners of code, data and models.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: Yes, we provide the details in section App.3.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: Yes, we do.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We only use it for writing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.